A coil of wire of area 0.2m^(2) containing 1000 turns is placed in a magnetic field of induction √2T. At time t=0 the axis of the coil and the direction of magnetic field are along z-direction.The coil is rotated by an angle 45° in 2s about y -axis. Assuming the constant angular speed the average e.m.f induced in the coil is approximately
Ans :41.4V
1
Expert's answer
2020-11-25T11:09:27-0500

Given quantities:

number of turns(n)=1000

Area of the coil(A)= 0.2 square meter

Magnetic field(B)= "\sqrt{2}" Tesla

Angle("\delta\theta") rotated by 45"\degree" in time("\delta{t}") 2 sec with constant angular velocity ("\omega") .


Detailed solution

formula for flux is...

"\phi=\vec{B}.\vec{A}= B A \cos{\theta}=B A \cos{\omega t}\\
emf=-\bigg(\cfrac{d\phi}{dt}\bigg)\\
emf=\omega n B A \sin{\omega t}\\
<emf>=\cfrac{\int_0^2{(emf) dt}}{\int_0^2{dt}}\\
<emf>=-\cfrac{nBA}{2}\bigg|\cos(\omega t)\bigg|_0^2\\
<emf>=-\cfrac{nBA}{2}[\cos{2\omega }- \cos{0}].....eq[1]\\"

and


"\delta\theta=\omega\delta t\\
\cfrac{\pi}{4}=\omega \times 2\\
\omega =\cfrac{\pi}{8}\\"

Now putting the value of "\omega" in eq[1] and we get...

"<emf>=-\cfrac{nBA}{2}\bigg[\cos(\cfrac{\pi}{4})-1\bigg]\\"

putting the values of n, B and A, we get..

"<emf>=\cfrac{1000\times \sqrt{2}\times0.2 }{2}\bigg[\cfrac{\sqrt{2}-1}{\sqrt{2}}\bigg]\\
=100(\sqrt2-1)\\
<emf>=41.4 volt.....Ans"

average e.m.f induced in the coil is 41.4 volt.